OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = −1

OpenStudy (anonymous):

@Michele_Laino @taramgrant0543664 @UsukiDoll

OpenStudy (exoticinspirit):

first What is the equation for a parabola

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

To find a quadratic equation from a parabola using distances, use the combined distance formula. \[\sqrt{(y_2-y_1)^2+(x_2-x_1)^2}=\sqrt{(y_2-y_1)^2}\] Plug in the focus on the left side of the equation. Plug in the directrix on the right side.

OpenStudy (triciaal):